OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!!! The table below shows the squares of different numbers: Number (x) 2 −2 3 −3 Square of the number (y) 4 4 9 9 Part A: Does the table represent y as a function of x? Justify your answer. (5 points) Part B: The total cost f(x), in dollars, for renting a rowboat for x hours is shown below: f(x) = 20 + 3x What is the value of f(150), and what does f(150) represent? (5 points)

OpenStudy (anonymous):

Will fan and medal for correct answer

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ParthKohli

OpenStudy (agl202):

A: I don't think that A is a function because 4 square rooted is 2 yes, but 4 square rooted isn't -2. the square of 9 is 3, not -3.

OpenStudy (agl202):

B: f(x) = 90 + 0.11x plug in 300 for x f(300) = 90 + 0.11(300) 90 is the cost of the car, 0.11 is the cost per mile so x would be any number he drove, it says to use 300, so it implies he drove 300 miles.

OpenStudy (anonymous):

why 300 hundred not 150

OpenStudy (agl202):

My bad let me redo that.

OpenStudy (anonymous):

@ParthKohli

OpenStudy (agl202):

Sorry for the wait(lag) Y is a function of x because every input value(x) will have only one outcome(y) f(150)=470 , it means that renting a rowboat for 150 hours will cost $450 +$20 for a flat fee

OpenStudy (anonymous):

what do you mean by a is correct i don't understand this question